# On the convergence of adaptive approximations for stochastic differential equations

### v4

@@ 1 Introduction
Stochastic differential equations (SDEs) have seen widespread use in the physical, engineering and social sciences for describing random systems. Examples of SDEs can be found within finance [3, 58, 67], biology [1, 4, 38, 72], physics [13, 45, 52, 68] and, more recently, in data science [8, 9, 12, 20, 35, 42, 48, 49, 70, 73]. However, just as for ordinary differential equations (ODEs), solutions of SDEs are rarely obtainable in closed form, and numerical methods are often required in practice.
In such applications, this is typically done through Monte Carlo simulation, where numerical solutions are computed using independently generated random variables (for example, corresponding to the increments of the underlying Brownian motion). By independently generating multiple numerical solutions, one can then use a Monte Carlo estimator to compute quantities relating to the average behaviour of the SDE. As motivation for this paper, we note that the status quo is to propagate numerical solutions of SDEs forwards in time using a fixed step size, whereas so-called “adaptive” step sizes have seen great success for discretizing ODEs (see [28, Section II.4] and [69]). We refer the reader to the book [30] for an accessible introduction to SDE numerics.
@@
In [24], Gaines and Lyons consider adaptive step size controls for SDE simulation where the current “candidate” step size $h_{k}$ at time $t_{k}$ may be halved depending on the increment of the underlying Brownian motion $W_{t_{k}+h_{k}}-W_{t_{k}}$ sampled on $t_{k},t_{k}+h_{k}]$. This approach is intuitive since it can allow the numerical method for (1) to adapt to the “large” fluctuations of Brownian motion which, despite having a low probability, can significantly increase the overall approximation error. However, such adaptive step sizes are clearly non-previsible and therefore carry theoretical and practical challenges.
@@
    In scientific machine learning, computing gradients of loss (or error) functions with respect to the parameters of a differential equation can be memory intensive [59]. As a consequence, memory-efficient extensions of the Brownian tree have been developed, such as the Virtual Brownian Tree [36, 48] and Brownian Interval [41].
@@
    It was show in [24, Theorem 4.3] that pathwise convergence in the adaptive step size setting does still hold provided that the numerical solution $Y=\{Y_{k}\}$ satisfies
@@
  • To motivate this pathwise convergence result, Gaines and Lyons present a simple counterexample [24, Section 4.1] demonstrating that the Euler-Maruyama method, which clearly does not satisfy the above condition, can converge to an incorrect (i.e. not Itô) solution of the SDE when the adaptive step sizes are non-previsible.
Therefore, it is proposed in [24] that second iterated integrals of Brownian motion should be finely discretized (with $o(h_{k})$ accuracy) to ensure the SDE solver converges. As a consequence, following [24], research has focused on either previsible step sizes [14, 31, 32, 39, 44, 53, 55, 63], non-previsible adaptive step sizes for commutative SDEs [5, 6, 19, 34, 36, 51, 60, 65] or adaptive high order schemes with iterated integrals [2].
Unlike previous works, we consider the use of non-previsible adaptive step sizes for simulating general SDEs – but where the solver does not require iterated integrals $\int_{t_{k}}^{t_{k+1}}\hskip-2.84526pt\big{(}W_{t}^{i}-W_{t_{k}}^{i}\big{)}\hskip 0.7113pt\circ\hskip 0.7113ptdW_{t}^{j}\hskip 0.7113pt$, which are computationally expensive to finely discretize [16].
@@
We say $\mathcal{D}$ is a no-skip dyadic partition if it is a no-skip partition and all of its times are of the form $k2^{-n}\cdot T$ where $n\geq 0$ and $0\leq k\leq 2^{n}$ are integers.
@@
The adaptive step sizes described in Gaines and Lyons [24] give no-skip partitions. In their case, the function $\Lambda_{l}$ will halve a step size depending on an estimate of the local error which is determined by the Brownian increments in $\mathcal{W}_{k}^{\hskip 0.7113ptl}\hskip 0.7113pt$. $L$ is the stopping time of when the proposed numerical solution is accepted at time $T$.
@@
    where $h_{k}$ is the step size used, $q\geq 0$ and the random vector $z_{k}$ is assumed to be independent from the Brownian motion. For example, $z_{k}$ could be the additional uniform random variable used within so-called “randomized” numerical methods [29, 33, 43, 54, 66]. In addition, we assume that each step of the method satisfies
@@
The condition (3) is much less restrictive than [24, Equation (4.6)] since it does not require the second iterated integrals of the Brownian motion to be finely discretized with $o(h)$ accuracy. In practice, (3) can be verified as, for example,
@@
where $H_{k}=\int_{t_{k}}^{t_{k+1}}\hskip-2.13394pt\big{(}\frac{1}{2}-\frac{t}{h_{k}}\big{)}\hskip 0.7113ptdW_{t}\sim\mathcal{N}(0,\frac{1}{12}h_{k}\hskip 0.7113ptI_{d})$ is independent of $W_{k}$ [18, Theorem 5.4]. Intuitively, the fact that these approximations are conditional expectations enables us to employ certain martingale arguments for establishing convergence in our analysis. Otherwise, the $O(h)$ local errors could propagate linearly to give a global $O(1)$ error.
@@ 1.1 Examples of adaptive numerical methods
where $H_{k}:=\int_{t_{k}}^{t_{k+1}}\hskip-0.7113pt\big{(}\frac{1}{2}-\frac{t}{h_{k}}\big{)}\hskip 0.7113ptdW_{t}=\frac{1}{h_{k}}\int_{t_{k}}^{t_{k+1}}\hskip-0.7113pt(W_{t}-W_{t_{k}}-\frac{t-t_{k}}{h_{k}}W_{k})\hskip 0.7113ptdt\sim\mathcal{N}(0,\frac{1}{12}h_{k}\hskip 0.7113ptI_{d})$ is the “space-time” Lévy area of Brownian motion on $t_{k}\hskip 0.7113pt,t_{k+1}]$, see [21, Definition 3.5].
@@
Whilst [24] showed that the Euler-Maruyama method may fail to converge to the Itô SDE solution when used with an adaptive step size, our main result establishes the convergence of the following “increment only” Milstein method introduced in [25].
@@
To our knowledge, the following are the first embedded stochastic Runge-Kutta methods for general SDEs (see [64, 65] for examples with scalar or additive noise). We first consider a multidimensional extension of the Heun scheme proposed in [62].
@@
In addition, we shall introduce a new stochastic Runge-Kutta method that uses space-time Lévy area $H_{k}$ to achieve a slightly better local accuracy than traditional “increment only” schemes, including Heun’s method (see Appendix A for details). For additive noise SDEs, the proposed scheme becomes high order and achieves an $O(h^{1.5})$ strong convergence rate (i.e. when the noise vector fields $g_{i}$ are constant). This scheme was inspired by a recent path-based splitting method [17, Example 1.3].
@@
In SDE numerics, it is common to consider mean squared errors. Informally speaking, in the constant step size regime, a local mean squared error of $O(h^{p})$ gives a global mean squared error of $O(h^{p-1})$ (see [52, Theorem 1.1] for details). Therefore, if we would like the global error to be less than $\varepsilon>0$, then we would expect local mean squared errors to be at most $O(\varepsilon h)$. This motivates the estimator (16), although $\|Y_{k+1}-\widetilde{Y}_{k+1}\|_{2}^{2}$ is a proxy for the pathwise (and not the mean squared) error.
@@ 1.2 Organisation of the paper
In Section 2, we will establish our pathwise convergence result (Theorem 2.26) for a wide class of adaptive SDEs approximations. The key idea in our proof methodology is to employ rough path theory and show that piecewise polynomial approximations of Brownian motion converge in the $\alpha\hskip 0.7113pt$-Hölder metric along any nested subsequence of the partitions produced by the SDE solver. This argument is a natural extension of the analysis presented in [23, Section 13.3.2], which presents similar results for piecewise linear approximations of Brownian motion that are defined along a sequence of partitions that are nested, but also deterministic.
Once we have $\alpha\hskip 0.7113pt$-Hölder convergence for the piecewise polynomial approximations, we can apply an extension of the Universal Limit Theorem from rough path theory [22] to show that the corresponding sequence of Controlled Differential Equations (CDEs) converges to the SDE solution almost surely. We show that pathwise convergence is still preserved when small $o(h)$ local perturbations are added to the CDE solutions. By considering the local Taylor expansion of the CDEs between discretization times, we obtain the condition (3) from Theorem 1.5 for the numerical method to satisfy. We conclude Section 2 by showing the no-area Milstein, Heun and Splitting Path Runge-Kutta (SPaRK) methods in Section 1.1 satisfy the convergence condition (3).
@@
In Section 4, we present an experiment comparing methods with different step size controls for the well-known SABR [^1] Stochstic Alpha-Beta-Rho model used in mathematical finance [7, 11, 27, 47]. The SABR model describes the evolution of an interest (or exchange) rate $S=\{S_{t}\}$ whose local volatility $\sigma=\{\sigma_{t}\}$ is also stochastic. It is given by the following Itô SDE:
@@
In addition, we also see that similar improved convergence rates can be obtained if local errors are estimated via an embedded Runge-Kutta method and subsequently used by a “Proportional-Integral” (PI) adaptive step size controller based on [6, 34]. However, due to our counterxample, we modify the PI step size controller so that it does not “skip over” times where Brownian information has previously been sampled. Based on the results of the experiment, we believe that our modified PI adaptive step size controller could be a promising technique for simulating non-commutative SDEs.
@@ 1.3 Notation
Throughout, $W=\{W_{t}\}_{t\in[0,T]}$ will be a standard $d$-dimensional Brownian motion. For a sequence of (random) no-skip partitions $\{D_{n}\}$ of $[0,T]$, we use $\widetilde{W}^{n}=\{\widetilde{W}_{t}^{n}\}_{t\in[0,T]}$ to denote a piecewise polynomial approximation of $W$ defined with respect to $\mathcal{D}_{n}\hskip 0.7113pt$. Most notably, this can be the usual piecewise linear discretization (see [21] for details). We write increments of the paths as $W_{s,t}:=W_{t}-W_{s}$ and $\widetilde{W}_{s,t}^{n}:=\widetilde{W}_{t}^{n}-\widetilde{W}_{s}^{n}$ for $s\leq t$.
We will denote SDE solutions by $y=\{y_{t}\}_{t\in[0,T]}$ and CDE solutions by $\widetilde{x}^{\hskip 0.7113ptn}$ or $\widetilde{y}^{\hskip 0.7113ptn}$. Each partition can be expressed as $\mathcal{D}_{n}=\{0=t_{0}^{n}<t_{1}^{n}<\cdots<t_{K_{n}-1}^{n}<t_{K_{n}}^{n}=T\}$ or $\{\mathcal{D}_{n}\}=\{h_{k}^{n}\}_{0\leq k<K_{n}}$ where $K_{n}\geq 1$ and $h_{k}^{n}:=t_{k+1}^{n}-t_{k}^{n}$ represents the step size. We often use the shorthand notation $W_{k}:=W_{t_{k+1}}-W_{t_{k}}$ when defining numerical methods – which are denoted by $Y=\{Y_{k}\}_{k\geq 0}$. Similarly, we write $H_{k}\sim\mathcal{N}(0,\frac{1}{12}h_{k}\hskip 0.7113ptI_{d})$ for the space-time Lévy area of Brownian motion on $t_{k}\hskip 0.7113pt,t_{k+1}]$, see [21, Definition 3.5].
@@ 2 Main result
We will first show the “rough path” convergence of piecewise polynomial approximations, defined on no-skip dyadic partitions, to Brownian motion. We note that, unlike [23, Section 13.3.2], which our analysis is based on, the sequence of nested partitions that we consider is allowed to depend on the Brownian motion. However, due to their no-skip property, our nested partitions will still produce a filtration – which will facilitate the martingale arguments we use to show convergence. We note that this strategy for showing convergence was originally proposed in the doctoral thesis [15, Chapter 6]. However, in this paper, we introduce the notion of no-skip dyadic partitions and consider a more general class of adaptive step sizes which can depend on random variables that are independent of the Brownian motion.

### v5

@@ 1 Introduction
Stochastic differential equations (SDEs) have seen widespread use in the physical, engineering and social sciences for describing random systems. Examples of SDEs can be found within finance [brigo2006finance, oksendal2003SDEs, shreve2004finance], biology [allen2010biology, browning2020biology, jha2012biology, vadillo2019LV], physics [strauss2017physics, leimkuhler2015ULD, milstein2004physics, sobczyk1991physics] and, more recently, in data science [chen2014SGHMC, cheng2018MCMC, dockhorn2022diffusion, foster2021shifted, salvi2023NSDEs, kidger2021NSDEs1, li2020NSDEs, li2019LangevinMC, song2021scoredbased, welling2011SGLD]. However, just as for ordinary differential equations (ODEs), solutions of SDEs are rarely obtainable in closed form, and numerical methods are often required in practice.
In such applications, this is typically done through Monte Carlo simulation, where numerical solutions are computed using independently generated random variables (for example, corresponding to the increments of the underlying Brownian motion). By independently generating multiple numerical solutions, one can then use a Monte Carlo estimator to compute quantities relating to the average behaviour of the SDE. As motivation for this paper, we note that the status quo is to propagate numerical solutions of SDEs forwards in time using a fixed step size, whereas so-called “adaptive” step sizes have seen great success for discretizing ODEs (see [hairer1993odes, Section II.4] and [soderlind2002odes]). We refer the reader to the book [higham2021SDEs] for an accessible introduction to SDE numerics.
@@
In [gaines1997variable], Gaines and Lyons consider adaptive step size controls for SDE simulation where the current “candidate” step size $h_{k}$ at time $t_{k}$ may be halved depending on the increment of the underlying Brownian motion $W_{t_{k}+h_{k}}-W_{t_{k}}$ sampled on $t_{k},t_{k}+h_{k}]$. This approach is intuitive since it can allow the numerical method for (1) to adapt to the “large” fluctuations of Brownian motion which, despite having a low probability, can significantly increase the overall approximation error. However, such adaptive step sizes are clearly non-previsible and therefore carry theoretical and practical challenges.
@@
    In scientific machine learning, computing gradients of loss (or error) functions with respect to the parameters of a differential equation can be memory intensive [rackauckas2021universal]. As a consequence, memory-efficient extensions of the Brownian tree have been developed, such as the Virtual Brownian Tree [jelincic2024VBT, li2020NSDEs] and Brownian Interval [kidger2021NSDEs2].
@@
    It was show in [gaines1997variable, Theorem 4.3] that pathwise convergence in the adaptive step size setting does still hold provided that the numerical solution $Y=\{Y_{k}\}$ satisfies
@@
  • To motivate this pathwise convergence result, Gaines and Lyons present a simple counterexample [gaines1997variable, Section 4.1] demonstrating that the Euler-Maruyama method, which clearly does not satisfy the above condition, can converge to an incorrect (i.e. not Itô) solution of the SDE when the adaptive step sizes are non-previsible.
Therefore, it is proposed in [gaines1997variable] that second iterated integrals of Brownian motion should be finely discretized (with $o(h_{k})$ accuracy) to ensure the SDE solver converges. As a consequence, following [gaines1997variable], research has focused on either previsible step sizes [fang2020adaptiveeuler, hoel2022highorderadaptive, hofmann2000additive, kelly2018adaptive, lamba2007adaptiveuler, mora2023variable, neuenkirch2019adaptiveeuler, romisch2006adaptive], non-previsible adaptive step sizes for commutative SDEs [burrage2002variable, burrage2004variable, foster2022sle, ilie2015adaptive, jelincic2024VBT, mauthner1998variable, rackauckas2017adaptive, adaptive2021brownian] or adaptive high order schemes with iterated integrals [bayer2023adaptive].
Unlike previous works, we consider the use of non-previsible adaptive step sizes for simulating general SDEs – but where the solver does not require iterated integrals $\int_{t_{k}}^{t_{k+1}}\hskip-2.84526pt\big(W_{t}^{i}-W_{t_{k}}^{i}\big)\hskip 0.7113pt\circ\hskip 0.7113ptdW_{t}^{j}\hskip 0.7113pt$, which are computationally expensive to finely discretize [foster2024signature].
@@
We say $\mathcal{D}$ is a no-skip dyadic partition of $[0,T]$ if it is a no-skip partition and, for each $\omega\in\Omega$, $\mathcal{D}(\omega)=\mathcal{D}_{N}(\omega)$ for some $N(\omega)\geq 0$ where $\mathcal{D}_{0}:=\{0,T\}$ and $\hskip 0.7113pt\mathcal{D}_{n+1}:=\{\frac{1}{2}(t_{k}+t_{k+1})\}\cup\mathcal{D}_{n}$ for some choice of consecutive times $t_{k}\hskip 0.7113pt,t_{k+1}\in\mathcal{D}_{n}\hskip 0.7113pt$.
@@
The adaptive step sizes described in Gaines and Lyons [gaines1997variable] give no-skip partitions. In their case, the function $\Lambda_{l}$ will halve a step size depending on an estimate of the local error which is determined by the Brownian increments in $\mathcal{W}_{k}^{\hskip 0.7113ptl}\hskip 0.7113pt$. $L$ is the stopping time of when the proposed numerical solution is accepted at time $T$.
@@
    where $h_{k}$ is the step size used, $q\geq 0$ and the random vector $z_{k}$ is assumed to be independent from the Brownian motion. For example, $z_{k}$ could be the additional uniform random variable used within so-called “randomized” numerical methods [he2020randomized, hu2019randomized, kruse2021randomized, morkisz2021randomized, shen2019randomized]. In addition, we assume that each step of the method satisfies
@@
The condition (3) is much less restrictive than [gaines1997variable, Equation (4.6)] since it does not require the second iterated integrals of the Brownian motion to be finely discretized with $o(h)$ accuracy. In practice, (3) can be verified as, for example,
@@
where $H_{k}=\int_{t_{k}}^{t_{k+1}}\hskip-2.13394pt\big(\frac{1}{2}-\frac{t}{h_{k}}\big)\hskip 0.7113ptdW_{t}\sim\mathcal{N}(0,\frac{1}{12}h_{k}\hskip 0.7113ptI_{d})$ is independent of $W_{k}$ [foster2023levyarea, Theorem 5.4]. Intuitively, the fact that these approximations are conditional expectations enables us to employ certain martingale arguments for establishing convergence in our analysis. Otherwise, the $O(h)$ local errors could propagate linearly to give a global $O(1)$ error.
@@ 1.1 Examples of adaptive numerical methods
where $H_{k}:=\int_{t_{k}}^{t_{k+1}}\hskip-0.7113pt\big(\frac{1}{2}-\frac{t}{h_{k}}\big)\hskip 0.7113ptdW_{t}=\frac{1}{h_{k}}\int_{t_{k}}^{t_{k+1}}\hskip-0.7113pt(W_{t}-W_{t_{k}}-\frac{t-t_{k}}{h_{k}}W_{k})\hskip 0.7113ptdt\sim\mathcal{N}(0,\frac{1}{12}h_{k}\hskip 0.7113ptI_{d})$ is the “space-time” Lévy area of Brownian motion on $t_{k}\hskip 0.7113pt,t_{k+1}]$, see [foster2020poly, Definition 3.5].
@@
Whilst [gaines1997variable] showed that the Euler-Maruyama method may fail to converge to the Itô SDE solution when used with an adaptive step size, our main result establishes the convergence of the following “increment only” Milstein method introduced in [giles2014antithetic].
@@
To our knowledge, the following are the first embedded stochastic Runge-Kutta methods for general SDEs (see [embed2003srk, adaptive2021brownian] for examples with scalar or additive noise). We first consider a multidimensional extension of the Heun scheme proposed in [roberts2012euler].
@@
In addition, we shall introduce a new stochastic Runge-Kutta method that uses space-time Lévy area $H_{k}$ to achieve a slightly better local accuracy than traditional “increment only” schemes, including Heun’s method (see Appendix A for details). For additive noise SDEs, the proposed scheme becomes high order and achieves an $O(h^{1.5})$ strong convergence rate (i.e. when the noise vector fields $g_{i}$ are constant). This scheme was inspired by a recent path-based splitting method [foster2024splitting, Example 1.3].
@@
In SDE numerics, it is common to consider mean squared errors. Informally speaking, in the constant step size regime, a local mean squared error of $O(h^{p})$ gives a global mean squared error of $O(h^{p-1})$ (see [milstein2004physics, Theorem 1.1] for details). Therefore, if we would like the global error to be less than $\varepsilon>0$, then we would expect local mean squared errors to be at most $O(\varepsilon h)$. This motivates the estimator (16), although $\|Y_{k+1}-\widetilde{Y}_{k+1}\|_{2}^{2}$ is a proxy for the pathwise (and not the mean squared) error.
@@ 1.2 Organisation of the paper
In Section 2, we will establish our pathwise convergence result (Theorem 2.26) for a wide class of adaptive SDEs approximations. The key idea in our proof methodology is to employ rough path theory and show that piecewise polynomial approximations of Brownian motion converge in the $\alpha\hskip 0.7113pt$-Hölder metric along any nested subsequence of the partitions produced by the SDE solver. This argument is a natural extension of the analysis presented in [frizvictoir2010roughpaths, Section 13.3.2], which presents similar results for piecewise linear approximations of Brownian motion that are defined along a sequence of partitions that are nested, but also deterministic.
Once we have $\alpha\hskip 0.7113pt$-Hölder convergence for the piecewise polynomial approximations, we can apply an extension of the Universal Limit Theorem from rough path theory [friz2024rsdes] to show that the corresponding sequence of Controlled Differential Equations (CDEs) converges to the SDE solution almost surely. We show that pathwise convergence is still preserved when small $o(h)$ local perturbations are added to the CDE solutions. By considering the local Taylor expansion of the CDEs between discretization times, we obtain the condition (3) from Theorem 1.5 for the numerical method to satisfy. We conclude Section 2 by showing the no-area Milstein, Heun and Splitting Path Runge-Kutta (SPaRK) methods in Section 1.1 satisfy the convergence condition (3).
@@
In Section 4, we present an experiment comparing methods with different step size controls for the well-known SABR [^1] Stochstic Alpha-Beta-Rho model used in mathematical finance [SABR2017Cai, SABR2018Cui, SABR2002Hagan, SABR2017Leitao]. The SABR model describes the evolution of an interest (or exchange) rate $S=\{S_{t}\}$ whose local volatility $\sigma=\{\sigma_{t}\}$ is also stochastic. It is given by the following Itô SDE:
@@
In addition, we also see that similar improved convergence rates can be obtained if local errors are estimated via an embedded Runge-Kutta method and subsequently used by a “Proportional-Integral” (PI) adaptive step size controller based on [burrage2004variable, ilie2015adaptive]. However, due to our counterxample, we modify the PI step size controller so that it does not “skip over” times where Brownian information has previously been sampled. Based on the results of the experiment, we believe that our modified PI adaptive step size controller could be a promising technique for simulating non-commutative SDEs.
@@ 1.3 Notation
Throughout, $W=\{W_{t}\}_{t\in[0,T]}$ will be a standard $d$-dimensional Brownian motion. For a sequence of (random) no-skip partitions $\{D_{n}\}$ of $[0,T]$, we use $\widetilde{W}^{n}=\{\widetilde{W}_{t}^{n}\}_{t\in[0,T]}$ to denote a piecewise polynomial approximation of $W$ defined with respect to $\mathcal{D}_{n}\hskip 0.7113pt$. Most notably, this can be the usual piecewise linear discretization (see [foster2020poly] for details). We write increments of the paths as $W_{s,t}:=W_{t}-W_{s}$ and $\widetilde{W}_{s,t}^{n}:=\widetilde{W}_{t}^{n}-\widetilde{W}_{s}^{n}$ for $s\leq t$.
We will denote SDE solutions by $y=\{y_{t}\}_{t\in[0,T]}$ and CDE solutions by $\widetilde{x}^{\hskip 0.7113ptn}$ or $\widetilde{y}^{\hskip 0.7113ptn}$. Each partition can be expressed as $\mathcal{D}_{n}=\{0=t_{0}^{n}<t_{1}^{n}<\cdots<t_{K_{n}-1}^{n}<t_{K_{n}}^{n}=T\}$ or $\{\mathcal{D}_{n}\}=\{h_{k}^{n}\}_{0\leq k<K_{n}}$ where $K_{n}\geq 1$ and $h_{k}^{n}:=t_{k+1}^{n}-t_{k}^{n}$ represents the step size. We often use the shorthand notation $W_{k}:=W_{t_{k+1}}-W_{t_{k}}$ when defining numerical methods – which are denoted by $Y=\{Y_{k}\}_{k\geq 0}$. Similarly, we write $H_{k}\sim\mathcal{N}(0,\frac{1}{12}h_{k}\hskip 0.7113ptI_{d})$ for the space-time Lévy area of Brownian motion on $t_{k}\hskip 0.7113pt,t_{k+1}]$, see [foster2020poly, Definition 3.5].
@@ 2 Main result
We will first show the “rough path” convergence of piecewise polynomial approximations, defined on no-skip dyadic partitions, to Brownian motion. We note that, unlike [frizvictoir2010roughpaths, Section 13.3.2], which our analysis is based on, the sequence of nested partitions that we consider is allowed to depend on the Brownian motion. However, due to their no-skip property, our nested partitions will still produce a filtration – which will facilitate the martingale arguments we use to show convergence. We note that this strategy for showing convergence was originally proposed in the doctoral thesis [foster2020thesis, Chapter 6]. However, in this paper, we introduce the notion of no-skip dyadic partitions and consider a more general class of adaptive step sizes which can depend on random variables that are independent of the Brownian motion.